The perimeter of a rectangle is 136m. The length is 8m more than twice the width. Find the dimensions?

Answer by checkley77(12844)

L=2W+8 FOR THE LENGTH.
W=THE WIDTH.
2(2W+8)+2W=136
4W+16+2W=136
6W=136-16
6W=120
W=120/6
W=20 FOR THE WIDTH
L=2*20+8=40+8=48 FOR THE LENGTH.
pROOF
2*48+2*20=136
96+40=136
136=136